given utility function u=x0.5y0.5 where px=12 birr py=4 birr and the income of the consumer is M=240 birr.

a)draw the budget line.

b)what happens to the original budget line if the budget falls by 25%

c)what happens to the original budget line if the price of x doubles

d)what happens to the original budget line if the price of y falls to 2 birr


1
Expert's answer
2022-01-30T13:49:18-0500

"U=(X^{0.5}Y^{0.5})"

"P_x= 12"

"P_y= 4"

m= 240

a) Budget line

12X+4Y=240



b) If the budget falls by 25%

New budget= 75% "\times 240= 180"

New budget line 12X+4Y=180

Diagramatically it can be shown as below


c) If price of X doubles

New price of X= 24

New budget line= 24X+4Y= 240

Diagramatically is as below;




d) If the Price of Y falls to 2

New budget line= 12X+2Y=240

The budget line moves upwards on the Y intercept as the consumer can afford more of Y while the X intercept remains constant. This can be explained diagramatically as below;